OpenStudy (anonymous):

What exponential function is the best fit for the data in the table? x f(x) 1 −4 3 −1 4 3

OpenStudy (anonymous):

a. f(x) = 4(3)x − 1 + 4 b. f(x) = 4(3)x − 1 − 4 c. f(x) = one fourth(3)x − 1 + 4 d. f(x) = one fourth(3)x − 1 − 4